How do I find y?
The problem is: x-2y=0 (equation 1) 3x+y=0 (equation 2)
Right now in my Algebra Honor class we're learning "Solving Systems by Substituion" and at first I understood it but now I can't seem to get the answer I want. I don't know what to do so please help!
Anna "Haylea"

Answer by usyim88hk(157)

OK, first we have to takes out both x and 3x from both equations, so lets multiply equation 1 by -3 so then we can takes the x-terms away from both equations:
:
(equation 1)
(-3)(x-2y) = 0(-3)
-3x+6y = 0
:
So now we can use substituion now
-3x+6y = 0
3x+y = 0
-----------
7y = 0
y = 0
--------------
So y is equal to zero